# Blending 3D Geometry and Machine Learning for Multi-View Stereopsis

### v1

@@ 5 Discussion
GC-MVSNet++ achieves new state-of-the-art on the DTU and BlendedMVS dataset. It achieves the second spot on the intermediate evaluation set and state-of-the-art on the advanced evaluation set of the Tanks and Temples dataset. We provide a short comparison with other recent methods.
GC-MVSNet++ vs.  MVSFormer
MVSFormer [54] focuses on using pre-trained models for feature extraction and extensive use of transformers. GC-MVSNet++ uses only one global feature aggregation module after feature extraction for global context enhancement. It focuses on effectively using geometric constraints across multiple views, and on improving the architecture of feature extraction and cost regularization networks.
GC-MVSNet++ vs. TransMVSNet
TransMVSNet [5] uses regular 2D convolution-based FPN (with batch-norm) for feature extraction and employs adaptive receptive field (ARF) modules with deformable layers after feature extraction. It trains using focal loss [63]. GC-MVSNet++ replaces the combination of regular FPN and ARF modules with weight-standardized deformable FPN (with group-norm) for feature extraction. It also uses a novel densely connected CostRegNet for cost volume regularization. It trains with a combination of cross-entropy loss and geometric consistency penalty for accelerated learning.
GC-MVSNet++ vs. CasMVSNet
CasMVSNet [10] proposes a coarse-to-fine regularization technique. It uses a regular 2D convolution-based FPN for feature extraction, generates variance-based cost volume, and employs depth regression to estimate $D_{0}$. The only similarity with our model is that we also use coarse-to-fine regularization.
@@ 6 Conclusion
In this paper, we introduce GC-MVSNet++, an enhanced learning-based MVS pipeline that explicitly models the geometric consistency of reference depth maps across multiple source views during training. Our approach incorporates a novel dense-CostRegNet for precise estimation of reference view depth maps. To our knowledge, this represents the first attempt to integrate multi-view, multi-scale geometric consistency checks into the training process. We demonstrate that the GC module is versatile and can be integrated with other MVS methods to enhance their learning efficiency. Through extensive experiments and ablation studies, we highlight the advantages of GC-MVSNet++. This work shows how to blend traditional geometric techniques and modern machine learning methods, to achieve more accurate and reliable 3D reconstructions from multiple images.

### v2

@@ 5 Discussion
We provide a comparison of our model with recent state-of-the-art models like MVSFormer, MVSFormer++, TransMVSNet and CasMVSNet.
GC-MVSNet++ vs. MVSFormer and MVSFormer++: MVSFormer focuses on using state-of-the-art pre-trained models for feature extraction and extensively use transformers in all its components. MVSFormer++ further adds multiple transformer modules in the MVS pipeline to leverage global feature aggregation in cost volume regularization. GC-MVSNet++ uses only one global feature aggregation module after feature extraction for global context enhancement before cost volume creation. It focuses on enforcing geometry-based learning across multiple views and on effective utilization of dense connections to enhance the cost regularization network.
GC-MVSNet++ vs. TransMVSNet: TransMVSNet uses regular 2D convolution-based FPN (with batch-norm) for feature extraction and employs adaptive receptive field (ARF) modules with deformable layers after feature extraction. It trains using focal loss [73]. GC-MVSNet++ replaces the combination of regular FPN and ARF modules with weight-standardized deformable FPN (with group-norm) for feature extraction. It also uses a novel densely connected CostRegNet for cost volume regularization. It trains with a combination of cross-entropy loss and geometric consistency penalty for accelerated learning.
GC-MVSNet++ vs. CasMVSNet: CasMVSNet [10] proposes a coarse-to-fine regularization technique. It uses a regular 2D convolution-based FPN for feature extraction, generates variance-based cost volume, and employs depth regression to estimate $D_{0}$. The only similarity with our model is that we also use coarse-to-fine regularization.
@@ 6 Conclusion
In this paper, we introduce GC-MVSNet++, an enhanced learning-based MVS pipeline that explicitly models the geometric consistency of reference depth maps across multiple source views during training. Our approach incorporates a novel dense-CostRegNet with two distinct modules, simple dense-module which is effective for encoding and feature-dense-module, which is designed to leverage dense connection for effective decoding and precise estimation of reference view depth maps. Ours is the first few method to integrate multi-view, multi-scale geometric consistency checks into the training process. We demonstrate that the GC module is versatile and can be integrated with other MVS methods to enhance their learning. Through extensive experiments and ablation studies, we highlight the advantages of GC-MVSNet++. This work shows how to blend traditional geometric techniques and modern machine learning methods, to achieve more accurate and reliable 3D reconstructions from multiple images.
Acknowledgement: This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the government of the Republic of Korea [25ZC1110, The research of the basic media $\cdot$ contents technologies].
Figure 8: The first row illustrates the selection of the $M$ closest source images for a given reference image. The middle row displays the corresponding ground truth depth maps, while the last row shows the remapped source depth maps, achieved by projecting the reference view’s x-y coordinates onto the source view. During remapping, any additional pixels from the source views are discarded. The remapped depths are then back-projected onto the source view to generate a mask. This reference view mask is applied to the per-pixel penalty to limit the penalties. The resulting final $\xi_{p}$ is presented in Fig. 3 of the paper. All depth maps are displayed within their respective view masks.
Appendix A Occlusion and its impact
Modeling pixel occlusion in multi-view settings presents a significant challenge, particularly when reasoning about pixels whose corresponding 3D points are occluded in other views. This issue becomes more pronounced when penalties are assigned to all such pixels, as in the proposed multi-view geometric consistency (GC) module. The GC module evaluates the geometric consistency of each pixel across multiple source views and imposes penalties for inconsistencies. However, penalizing occluded pixels and combining these penalties with depth errors negatively affects the training process. In our initial experiments, we observed that this approach caused the loss to explode during training, meaning the loss values increased as the model continued to train.
Our investigation reveals that incorrect penalties applied to occluded pixels dominated the loss during training. To address this, we implemented a series of steps that made our method more robust to this issue. First, we selected the closest source view images, as defined by MVSNet [15], to reduce the likelihood of occluded pixels. The first row of Fig. 8 illustrates this source view selection for a given reference view. Using the nearest views minimizes the number of potentially occluded pixels.
Second, during forward-backward reprojection, we remap the source view depth map based on the x-y coordinate projections from the reference view to the source view, and then back-project these remapped values to the reference view (as described in Alg. 2 of the paper). The last row of Fig. 8 shows the remapped source view depth maps. During this remapping process, occluded and extraneous pixels from the source view are discarded, ensuring that only valid pixels are back-projected. This step effectively manages extreme cases of occlusion and additional visible pixels.
Finally, after generating the per-pixel penalty, we apply a binary mask from the reference view to exclude any pixels not part of the scene (see Fig. 3 in the paper). This combination of steps significantly reduces the impact of incorrect penalties and stabilizes the training process.
Appendix B Geometric Consistency Module
Figure 9: GC module flow-chart for consistency check.
We describe the steps of the geometric consistency (GC) module in Fig. 9. At each stage, the geometric consistency of the estimated depth map is checked across $M$ source views. For each source view, we perform the forward-backward-reprojection of estimated depth map to reason about geometric inconsistency of pixels (described in Alg. 2). In this three-step process, first, we warp each pixel $P_{0}$ of a reference view depth map $D_{0}$ to its $i^{th}$ neighboring source view to obtain corresponding pixel $P^{{}^{\prime}}_{i}$. Then, we back-project $P^{{}^{\prime}}_{i}$ into 3D space and finally, we reproject it to the reference view as $P^{"}_{0}$ using $c_{0}$. $D_{0}$, $D^{{}^{\prime}}_{P^{{}^{\prime}}_{i}}$ and $D^{"}_{P^{"}_{0}}$ represents depth value of pixels associated with $P_{0}$, $P^{{}^{\prime}}_{i}$ and $P^{"}_{0}$ [49]. With $P^{"}_{0}$ and $D^{"}_{P^{"}_{0}}$, we calculate pixel displacement error (PDE) and relative depth difference (RDD). After taking logical-OR between PDE and RDD, we assign value $1$ to all inconsistent pixels and zero to all other pixels. The geometric inconsistency mask sum is generated over $M$ source views and averaged to generate per-pixel penalty $\xi_{p}$.
Appendix C Depth Interval Ratio (DIR)
Table 8: The performance of GC-MVSNet on evaluation set of DTU [21] with change in stage-wise DIR (depth interval ratio).
[…]
$[1,3]$ 2.0, 0.8, 0.40 0.338 0.269 0.3035
$[1,3]$ 2.0, 0.7, 0.35 0.343 0.264 0.3035
$[1,3]$ 2.0, 0.7, 0.30 0.331 0.27 0.3005
$[1,3]$ 1.6, 0.7, 0.30 0.329 0.271 0.300
DIR directly impacts the separation of two hypothesis planes at pixel level. For a given stage, the pixel-level depth interval is calculated as product of $DIR_{stage}$ and depth interval (DI). The value of DI is calculated using interval scale and a constant value provided in DTU camera parameter files.
Following the trend of modern learning-based methods [14, 10, 15, 16, 5, 6, 12], we train our model on $512\times 640$ resolution and test on $864\times 1152$ resolution. To adjust for the pixel-level depth interval caused by the increase in resolution, we explore different DIR values for testing on DTU. We train our model with stage-wise DIR $2.0,0.8,0.4$ ($DIR_{train}$). such that the refine stage pixel-level depth interval is same as the provided interval scale value of $1.06$. Table 8 shows DIR values for evaluation on DTU, we only explore smaller values than $DIR_{train}$ to compensate for the increase in resolution. GC-MVSNet achieves its optimal performance at DIR $1.6,0.7,0.3$ with $\xi_{p}\in[1,3]$, $DIR_{test}$. We use the same $DIR_{train}$ and $DIR_{test}$ with $\xi_{p}\in[1,2]$.
Appendix D Depth Map Fusion Methods
The quality of point clouds depends heavily on depth fusion methods and their hyperparameters. Following the recent learning-based methods [5, 6, 10], we also use different fusion methods for DTU and Tanks and Temples dataset. For DTU, we use Fusibile [1] and for Tanks and Temples, we use Dynamic method [5, 13].
The fusibile fusion method uses three hyperparameters, disparity threshold, probability confidence threshold, and consistency threshold. The disparity threshold defines the upper limit of disparity for points to be eligible for fusion. The probability confidence threshold defines the lower limit of confidence above which points are eligible for fusion. The consistency threshold mandates that the eligible points be geometrically consistent across as many source views. During the fusion process, only those points that satisfy all three conditions are fused into a point cloud.
The dynamic fusion method uses only two hyperparameters, probability confidence threshold and consistency threshold. Both these hyperparameters have exact same function as in the Fusibile method. The disparity threshold is not provided by the user, it is dynamically adjusted during the fusion process.
Appendix E Accuracy and Completeness Metrics
Figure 10: The process of calculating accuracy and completeness for DTU [21] point cloud evaluation.
Accuracy and completeness are two metrics used with the DTU [21] dataset. Fig. 10 shows the process of calculation. Accuracy is the average of the distance of the first neighbor from the predicted point cloud to the ground truth point cloud. It only considers the points which are below the maximum threshold for the distance. For completeness, the same process is repeated but with ground truth as a referenced point cloud, i.e. the average of the distance of the first neighbor from the ground truth point cloud to the predicted point cloud.
Appendix F Inference Memory Comparison
Table 9: Illustration of model memory during the inference phase of various models at $864\times 1152$ and $1152\times 1536$ resolutions. We use information from MVSFormer++ paper to fill out CasMVSNet and TransMVSNet information.
Methods Resolution Depth Interval Memory (MB)
MVSFormer++ 864 $\times$1152 32-16-8-4 4873
64-32-8-4 5025
1152 $\times$1536 32-16-8-4 5964
64-32-8-4 6753
CasMVSNet 864 $\times$1152 48-32-8 4769
1152 $\times$ 1536 6672
TransMVSNet 864 $\times$ 1152 48-32-8 3429
1152 $\times$1536 6320
GC-MVSNet 864 $\times$1152 48-32-8 2787
1152 $\times$ 1536 4831
GC-MVSNet++ 864 $\times$1152 48-32-8 5221
1152 $\times$ 1536 8193
The memory consumption analysis in Table 9 demonstrates that while our proposed GC-MVSNet++ incurs increased memory usage compared to lighter architectures like GC-MVSNet, it remains competitive with transformer-based state-of-the-art methods such as MVSFormer++. This additional memory overhead results from the incorporation of dense connectivity in the cost volume regularization module, which significantly enhances representational flexibility and multi-scale information propagation. Crucially, this design choice leads to competitive depth estimation accuracy and robust geometric consistency, clearly justifying the increased memory footprint. Our model achieves a favorable balance, offering accuracy comparable to transformer-based methods without the added complexity and overhead of transformer layers.
Appendix G Use of Existing Assets
We use PyTorch to implement GC-MVSNet. It is based on CasMVSNet [10] and TransMVSNet [5]. These two methods heavily borrow code from the PyTorch implementation of MVSNet [15].
We use preprocessed images and camera parameters of DTU [21] dataset from the official repository of MVSNet [15] and R-MVSNet [8]. We follow [64] for training and testing on BlendedMVS [20]. For Tanks and Temples [22] evaluation, we use images and camera parameters as used in R-MVSNet [8].
Appendix H $e_{1}$ Error Comparison on BlendedMVS
In this section, we discuss the $e_{1}$-error plots for the GC‑MVSNet++ and TransMVSNet methods. A quantitative comparison table is provided in the main paper. We argue that training MVS methods with geometry‑based guidance not only accelerates the optimization process but also significantly improves depth‑estimation consistency. Figure 11 shows the error-map comparison with TransMVSNet. The first column displays images from a BlendedMVS test‑set scenes; the second and third columns show the absolute-error plots (darker red is higher error) for GC‑MVSNet++ and TransMVSNet, respectively. Although there are visual differences in these error maps, the fourth column reveals more detail: it presents a binary map (0 or 1) of all points where GC‑MVSNet++’s error is lower than TransMVSNet’s. This clearly demonstrates the advantage of geometry‑guided training over non‑geometry‑guided methods in MVS. While most test‑scene images exhibit visible improvements in their error-maps, these gains do not translate linearly into the final point‑cloud quality.
Figure 11: EPE error comparison of GC-MVSNet++ with TransMVSNet on validation set of BlendedMVS dataset. First column shows the image of the scene, second and third columns show their respective EPE error maps (darker is higher error), and last column shows the binary map (0,1) of all the points where GC-MVSNet++ has smaller EPE-error as compared to TransMVSNet.
Appendix I Point Clouds
In this section, we show all evaluation set points clouds reconstructed using GC-MVSNet on DTU [21], Tanks and Temples [22] and BlendedMVS [20] datasets. Fig. 12, 13 and 14 show all evaluation set point clouds from DTU, Tanks and Temples and BlendedMVS, respectively.
Figure 12: Point clouds reconstructed using GC-MVSNet for all scenes from DTU [21] evaluation set.
Figure 13: Point clouds reconstructed using GC-MVSNet for all scenes from Tanks and Temples [22] intermediate and advanced set.
Figure 14: Point clouds reconstructed using GC-MVSNet for all scenes from BlendedMVS [20] evaluation set.